OpenStudy (samantha2234):

What is the slope of the line passing through the points (0, -5) and (4, 2)?

OpenStudy (samantha2234):

@shansha1

OpenStudy (samantha2234):

Also what is the slope 3x+4y=12

OpenStudy (smilie2015):

slope = rise/run =\[\frac{ y2-y1 }{ x2-x1 }\]

OpenStudy (samantha2234):

So.... what does that mean?

OpenStudy (smilie2015):

so here we have (0, -5) and (4, 2 ) (0,-5) =(x1,y1) (4,2)=(x2,y2) put these values in the eq you will get the answer

OpenStudy (smilie2015):

try to do if there is still any problem let me know :)

OpenStudy (samantha2234):

Do I have to graph it? I really suck at graphing!!!

OpenStudy (smilie2015):

no it is not required in the question just solve it

OpenStudy (samantha2234):

I have no idea what to do

OpenStudy (smilie2015):

(0,-5) =(x1,y1) (4,2)=(x2,y2) put these values in the following equation : \[\frac{ y2-y1 }{ x2-x1 }\]

OpenStudy (samantha2234):

So I have to figure out what y2 and so on equals?

OpenStudy (smilie2015):

you have y2 just put the values for example (1,2) (3,4) here (x1,y1) are (1,2) and (x2,y2) are (3,4) formula is \[\frac{ y2-y1 }{ x2-x1}\] putting the values \[\frac{ 4-2 }{ 3-1 }\] 2/2=1

OpenStudy (samantha2234):

so you would substitute and then simplify?

OpenStudy (smilie2015):

yup :)

OpenStudy (samantha2234):

:) Thanks so much! and 0 would = y1 and -5 would = x1 4 = x2 and 2 =y2?

OpenStudy (smilie2015):

x1= 0 y1=-5 x2=4 y2=2

OpenStudy (smilie2015):

my pleasure :)

OpenStudy (samantha2234):

@smilie2015 is a lifesaver!!!

OpenStudy (smilie2015):

thnx

OpenStudy (samantha2234):

:) Have a nice day!